# ANUBHUTI: A Comprehensive Corpus for Sentiment Analysis in Bangla Regional Languages

### v1

@@ 3.5.2 Missing Data Detection
To ensure the completeness and reliability of the translated dataset, a dual-phase approach was adopted for identifying missing or incomplete entries. This involved both manual inspection and automated detection using Python scripting.
@@
Each translated entry was validated against its original Bangla source to ensure semantic completeness and fidelity of sentiment expression, thereby reducing the risk of information loss during translation.
@@
Algorithm 1 Check for Missing Data in CSV File
@@ 3.5.3 Anomaly Detection
In addition to manual checking, a semi-automated Python-based anomaly detection script was developed to systematically flag problematic entries. The script operated according to the following step, where the system is evaluated sentence formatting to identify unusual patterns such as excessive punctuation (e.g., “!!!!”, “….”), non-Bangla characters or scripts, and other spam-like structures. Such anomalies were considered indicators of noise or stylistic inconsistency that could impact downstream analysis.
@@
Algorithm 2 Detect Anomalies in Bangla Sentiment Dataset
@@ 3.5.4 Spelling Verification
Given the informal nature of written regional dialects, this step was essential in preserving clarity, authenticity, and interpretability in the translated texts.
The manual inconsistency checking process played a foundational role in enhancing the reliability and validity of the ANUBHUTI dataset. By ensuring translation accuracy, completeness, contextual consistency, and dialect-specific spelling correctness, the dataset was rendered suitable for training robust and dialect-aware sentiment analysis models. This quality control phase ensures that the models can effectively interpret and classify emotional content in diverse forms of Bangla, particularly those originating from underrepresented regional communities.
@@ 3.6 Data Annotation
To prepare the ANUBHUTI dataset for supervised sentiment analysis, a two-level annotation process was carried out. The annotation involved manual labeling of each sentence for both its thematic class (multiclass) and emotional sentiment (multilabel). These annotations enable the development and evaluation of models capable of understanding complex emotional expression and contextual bias within Bangla regional dialects. The decision to implement both multiclass and multilabel annotation schemes was motivated by the need to capture the full complexity of sentiment expression in regional Bangla dialects. The multiclass annotation provides a high-level thematic categorization—distinguishing whether a sentence is political, religious, or neutral—allowing for context-aware analysis of sentiment trends. Meanwhile, the multilabel emotion tagging captures the diverse and overlapping emotional states conveyed in informal dialectal speech.
3.6.1 Multiclass Annotation
@@ 3.6.2 Multilabel Annotation
3.6.3 Annotation Execution
The annotation of the ANUBHUTI dataset was carried out manually by the same regional dialect experts who were previously involved in the translation of the dataset. Given their linguistic familiarity and contextual awareness, these translators were well-positioned to interpret both the literal meaning and the emotional undertone of the texts within their dialectal context.
@@ 3.6.3 Annotation Execution
Table 4: Overview of Dataset Annotation for Chittagong region
@@ 3.7 Dataset Statistics
The ANUBHUTI dataset comprises a total of 2,000 sentences distributed equally across four major regional dialects of Bangla. As shown in Table 5, each region—Mymensingh, Noakhali, Sylhet, and Chittagong—contributes 500 sentences to ensure a balanced and comparable representation of dialectal variations. This equal allocation supports robust dialect-specific analysis and allows for fair model training and evaluation across regions. The dataset’s regional symmetry was a key design choice to avoid dialectal bias and to ensure methodological consistency in downstream sentiment analysis tasks.
@@
Mymensingh 500
Noakhali 500
Sylhet 500
Chittagong 500
Total 2000
Figure 2 presents the thematic classification of sentences into three mutually exclusive categories: Political, Religious, and Neutral. This distribution highlights the predominance of political content in the dataset, reflecting the socio-political context prevalent in the source material.
Figure 2: Multiclass Distribution Of Anubhuti Dataset
Figure 3 depicts the frequency of emotion labels assigned in the multilabel annotation scheme. The chart reveals that emotions such as Anger, Contempt, and Disgust occur more frequently, indicating the emotionally charged nature of the regional dialect texts, especially in political and religious discussions. Conversely, less frequent labels like Fear and Surprise reflect more nuanced or less commonly expressed sentiments.
Figure 3: Multilabel Distribution Of Anubhuti Dataset
@@
3.8 Data Distribution
The ANUBHUTI dataset was constructed to ensure balanced and representative coverage across multiple dimensions, including regional dialects, emotional sentiment categories, contextual themes, and machine learning data splits. A total of 2,000 dialectal sentences were used in the initial version of the dataset, with 500 sentences allocated per dialect region: Mymensingh, Noakhali, Sylhet, and Chittagong.
Dialectal Distribution
The dataset maintains an equal distribution across the four selected Bangla regional dialects. This facilitates fair comparison in dialect-specific sentiment analysis and ensures that models trained on the dataset are exposed to a balanced variety of regional linguistic patterns.
Sentiment and Contextual Label Distribution
Each sentence was manually annotated using a multilabel sentiment schema comprising seven emotional classes: anger, surprise, contempt, disgust, fear, enjoyment, and sadness. Additionally, a single contextual label was assigned to each sentence in a mutually exclusive fashion: political, religious, or neutral.
This dual-layer annotation allows for comprehensive modeling of both affective and topical dimensions. Care was taken to avoid extreme class imbalance across these categories, thereby supporting reliable supervised learning.
Train-Validation-Test Split
To prepare the dataset for machine learning experiments, the entire annotated dataset was randomly partitioned into three subsets:
  • Training Set (70%): Used to train sentiment classification models.
  • Validation Set (15%): Used to fine-tune model parameters and prevent overfitting.
  • Test Set (15%): Used to evaluate the final model’s generalization performance on unseen data.
Stratified sampling was applied during the splitting process to preserve the proportional distribution of sentiment labels and context classes across all three subsets. This approach ensures that each split contains a representative sample of the emotional and thematic diversity present in the dataset, which is critical for training robust and unbiased models.
4 Limitations
@@ 5 Ethics Statements
The data used to construct the ANUBHUTI dataset does not raise any ethical concerns, as it was collected from a publicly available dataset and through Manual Translation. The dataset contains no sensitive or private information, ensuring compliance with ethical standards. Additionally, no data was sourced from personal communications or restricted sources. This dataset is intended solely for NLP research and development, and no human or animal subjects were involved in its creation.
6 Data Availability
The ANUBHUTI dataset, which contains annotated sentiment analysis data in Bangla regional dialects, is publicly available for research and academic purposes. Researchers interested in utilizing the dataset can access it through the following link:Dataset of Sentiment Analysis for Regional Bangla Language (Original data) (Mendeley Data)
@@ 7 CRediT Author Statement
8 Acknowledgements
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
9 Declaration of Competing Interest

### v2

@@ 3.5.2 Missing Data Detection
To ensure completeness and reliability, we used a dual-phase approach to identify missing or incomplete translations. This involved both manual inspection and automated detection using Python scripting.
@@
Each translated entry was validated against its original Bangla source to ensure semantic completeness and fidelity of sentiment expression, reducing the risk of information loss during translation.
@@
Algorithm 1 Missing data check results in the ANUBHUTI dataset CSV file, displaying the count and percentage of missing values for each column
@@ 3.5.3 Anomaly Detection
In addition to manual checking, we developed a semi-automated Python script to systematically flag anomalous entries. The script evaluated sentence formatting to identify unusual patterns such as excessive punctuation (e.g., “!!!!”, “….”), non-Bangla characters or scripts, and other spam-like structures. Such anomalies were considered indicators of noise or stylistic inconsistency that could impact downstream analysis.
@@
Algorithm 2 Anomaly detection results in the Bangla sentiment dataset, highlighting data points identified as outliers
@@ 3.5.4 Dialectal Spelling Standardization
Given the informal nature of written regional dialects, this step was essential for preserving clarity, authenticity, and interpretability in the translated texts.
The manual inconsistency checking process played a foundational role in enhancing ANUBHUTI’s reliability and validity. By ensuring translation accuracy, completeness, contextual consistency, and dialect-specific spelling correctness, the dataset was rendered suitable for training robust and dialect-aware sentiment analysis models. This quality control phase ensures that models can effectively interpret and classify emotional content in diverse forms of Bangla, particularly from underrepresented regional communities.
@@ 3.6 Data Annotation
To prepare the ANUBHUTI dataset for supervised sentiment analysis, a two-level annotation process was carried out. The annotation involved manually labeling each sentence for its thematic class (multiclass) and emotional sentiment (multilabel). These annotations enable developing and evaluating models capable of understanding complex emotional expression and contextual bias in Bangla’s regional dialects. The decision to implement both multiclass and multilabel annotation schemes was motivated by the need to capture the full complexity of sentiment expression in regional Bangla dialects. The multiclass annotation provides a high-level thematic categorization—distinguishing whether a sentence is political, religious, or neutral—allowing for context-aware analysis of sentiment trends. Meanwhile, the multilabel emotion tagging captures the diverse and overlapping emotional states conveyed in informal dialectal speech.
3.6.1 Thematic Label Assignment
@@ 3.6.2 Emotional Label Assignment
3.6.3 Annotation Process
The annotation of the ANUBHUTI dataset was carried out manually by the same regional dialect experts who were previously involved in the translation of the dataset. Their linguistic familiarity and contextual awareness made these translators well-positioned to interpret the texts’ literal meaning and emotional undertone within their dialectal context.
@@ 3.6.3 Annotation Process
Table 3: Overview of dataset annotation process for the Chittagong region, detailing the annotation guidelines, tools used, and quality control measures
@@
Figure 2: Multiclass distribution of the ANUBHUTI dataset, showing the number of samples for each class label
@@ 3.7 Dataset Statistics
Mymensingh 2500
Noakhali 2500
Sylhet 2500
Chittagong 2500
Total 10,000
The ANUBHUTI dataset comprises a total of 10,000 sentences distributed equally across four major regional dialects of Bangla. As shown in Table 4, each region—Mymensingh, Noakhali, Sylhet, and Chittagong—contributes 2500 sentences to ensure a balanced and comparable representation of dialectal variations. This equal allocation supports robust dialect-specific analysis and allows for fair model training and evaluation across regions. The dataset’s regional symmetry was a key design choice to avoid dialectal bias and to ensure methodological consistency in downstream sentiment analysis tasks.
Figure 2 presents the thematic classification of sentences into three mutually exclusive categories: Political, Religious, and Neutral. Figure 3 depicts the frequency of emotion labels assigned in the multilabel annotation scheme. The chart reveals that emotions such as Anger, Contempt, and Disgust occur more frequently, indicating the emotionally charged nature of the regional dialect texts, especially in political and religious discussions. Conversely, less frequent labels like Fear and Surprise reflect more nuanced or less commonly expressed sentiments.
@@
Figure 3: Multilabel distribution of the ANUBHUTI dataset, displaying the frequency of each label combination within the dataset
4 Limitations and Future Works
@@
5 Conclusion
In this study, we introduced ANUBHUTI, the first comprehensive sentiment analysis corpus tailored to Bangla regional dialects—namely Mymensingh, Noakhali, Sylhet, and Chittagong. By manually translating and annotating 2,500 dialect-specific sentences with both thematic and emotional labels, we addressed a critical gap in low-resource NLP for Bangla. The dual-layer annotation schema—combining multiclass contextual categorization and multilabel emotional sentiment—enables a richer, more nuanced understanding of public discourse, especially in politically and religiously sensitive domains. Our methodical approach to data quality, including native translator involvement, Cohen’s Kappa-based agreement validation, and multi-stage inconsistency checks, ensures that ANUBHUTI stands as a reliable resource for dialect-aware sentiment modeling. The balanced representation across dialects and the inclusion of diverse emotions empower researchers to develop robust, culturally grounded NLP systems capable of interpreting underrepresented linguistic communities.
Looking ahead, ANUBHUTI lays the foundation for building sentiment-aware chatbots, misinformation detectors, and mental health assessment tools in regional Bangla. We envision future expansions that include more dialects, broader topical coverage, and real-world applications. By releasing this dataset publicly, we aim to catalyze inclusive, dialect-sensitive NLP research and promote linguistic equity within Bangla language technology development.
@@ 6 Ethics Statements
The data used to construct ANUBHUTI does not raise ethical concerns, as it was collected from a publicly available dataset and through manual translation. The dataset contains no sensitive or private information, ensuring compliance with ethical standards. Additionally, no data was sourced from personal communications or restricted sources. This dataset is intended solely for NLP research and development; no human or animal subjects were involved in its creation.
7 Availability of ANUBHUTI
The ANUBHUTI dataset, containing annotated sentiment analysis data in Bangla regional dialects, is publicly available for research and academic purposes. Researchers interested in utilizing the dataset can access it through the following link:Dataset of Sentiment Analysis for Regional Bangla Language (Original data) (Mendeley Data)
@@ 8 CRediT Author Statement
9 Competing Interests Statement